OpenStudy (anonymous):

medal For Help ASAP :) PLEAAASSSE The password for a school's grading system is a 4 digit code. If the password is composed of digits one through nine.... What is the probability that the password is composed of a one, a two, a three, and a four? A.1/5,040 B.4/9 C.1/6,561 D.1/36 I think it's D ? Am I correct ?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@mtbender74 @Luigi0210

OpenStudy (anonymous):

Can you digits repeat in the password?

OpenStudy (anonymous):

i would assume, yes...

OpenStudy (anonymous):

Just add up all the probabilities of each number being used. Like this The chance of it being a 1 is 1/9 The chance of it being a 2 is 1/9 So far I have 2/9.. Just add them all up

OpenStudy (anonymous):

So the number of codes that have 1, 2, 3, and 4 in any order would be 4*3*2*1 The total number of codes possible would be 9*9*9*9... first number is numerator, second is denominator...

OpenStudy (anonymous):

You have a total of 9^4 passwords possible. Out of these, the passwords using the digits 1, 2, 3 and 4 only are 4*3*2*1 = 24. Hence probability of getting a password composed of the digits 1,2,3 and 4 is 24/9^4

OpenStudy (anonymous):

Oh ok so it's b ?- I mulit them

OpenStudy (anonymous):

From the answer choices it seems that the question is asking the probability of finding the password 1234 and not any password with these digits in any order. In this case, the probability is 1 over 9^4 = 1/6561 or choice C

OpenStudy (anonymous):

@DemolisionWolf C or B ?

OpenStudy (anonymous):

@mathslover c or B

OpenStudy (anonymous):

If that is true @navk, i like this question less than when i first saw it :) But i see what you mean...i hate poorly written questions.

OpenStudy (anonymous):

@mtbender74 can it be B ?

mathslover (mathslover):

I think it is C! \(\cfrac{1}{9^4}\)

OpenStudy (anonymous):

@navk i don't think so...i looked at that but that would include passcodes 1111, 2222, etc. The problem says that the passcode has a 1, 2, 3, *and* 4...if that "and" said "or" i would agree with B

OpenStudy (anonymous):

Then it's C for sure